in a two dimensional motion of a particle, the particle moves from point A, position vector r1 to point B position vector r2. if the magnitude of these vector are respectively r1=3 and r2=4 and the angle they make with the x-axis are 75° and 15° respectively, then magnitude of the displacement vector is

Grade:11

1 Answers

Aman gohel
6 years ago
The angle between the vectors=75-15=60°=>Displacement=|r2-r1|=r2+(-r1)=>=√(r1^2+(-r2)^2+2r1r2cosX)=>√(9+16+24cos60°)=>√(9+16+12)=>√37=6.08m(Ans)